# ReactionMamba: Generating Short & Long Human Reaction Sequences

### v1

@@ IV-B Evaluation Metrics
A lower FID indicates that the generated data distribution closely matches the real data distribution, reflecting higher quality and diversity in the generated motions  [17, 25].
@@
where $S^{i}$ and $S^{j}$ are two sampled motion sequences from set $S$ (either real or generated), and $P$ is the set of randomly selected unique pairs. A diversity score close to the diversity score of the ground truth reflects greater variability within the motion set and a preservation of the original diversity.
IV-C Baselines
We compare our method with three SOTA baselines, ReMoS [20], Interformer [9] both retrained on NVIDIA H100 SXM5 GPU (80 GB HBM3 VRAM) and Ready-to-React [2] trained on NVIDIA Quadro RTX 6000 GPU (24 GB VRAM). ReMoS [20] framework consists of two modules a body diffusion module and a hand diffusion module that need to be trained separately. We retrained ReMos on each of the 3 datasets used in the experiments following instructions and recommendations from the authors github repository [19].
Interformer [9] is a Transformer-based model for reaction generation that we retrained on each of the 3 datasets used in the experiments following the protocol used by [20]. Ready-to-react [2], in contrast, employs an online reaction policy that conditions motion generation on the past trajectories of both interacting agents, enabling immediate and context-aware responses. In our experiments, we specifically adopt the reaction generation mode of Ready-to-React, which allows to generate the agents future movement based on its past history and the opponents motion. However, due to issues with preprocessing while following the protocol of the original implementation [66], we do not present results of Ready-to-React on the InterX dataset for the sake of fairness.
IV-D Implementation
Reaction Mamba was trained using Adam optimizer. The base learning rate is set to $10^{-4}$, it decays following a Cosine Annealing LR scheduler. The input dimension is set to $d=k\times 3$, $d_{model}=256$, $d_{z}=128,d_{intermediate}=512$. We use 6 layers of the mamba block for both encoder and decoder. Before concatenation, the action and initial frame are projected to $d_{c}=64$ each of them. We set the loss weights, $w_{\mathrm{recon}}=w_{\mathrm{recon}}=1$ and $w_{\mathrm{KL}}=0.5$. Reaction Mamba was trained on NVIDIA Quadro RTX 6000 GPU (24 GB VRAM), for 300k iterations for Lindy Hop and 90K iterations for NTU120-AS and InterX.
V Results
Quantitative Results. In Table I we present results on LindyHop, and NTU120-AS, using each dataset default sequence length. We observe that ReactionMamba is competitive with the state-of-the-art on LindyHop across all metrics. Ready-to-React achieves the best MPJPE and MPJVE on these shorter sequences, showing its strength in short-horizon positional accuracy. However, its FID scores are considerably higher, revealing a lack of realism and consistency in longer rollouts. Meanwhile, we outperform all baselines on NTU120-AS, which can be explained by the lack of explicit contact information in this dataset. Indeed, this information is used by ReMoS to increase generation accuracy but is not needed by ReactionMamba. Another reason is ReactionMamba’s ability to better deal with sequences where the reactor does not move away from its original position (e.g. a sequence where two persons stand and shake hands). Our model is designed to inject the initial pose information at each frame in the conditioning module. Since the NTU120-AS dataset contains many such sequences, ReactionMamba is able to outperform InterFormer and Ready-to-React.
In Table II, we present results for long-term generation on the InterX dataset with 200-frame-long sequences. ReactionMamba clearly outperforms all baselines across metrics, generating accurate, smooth, and diverse long-horizon sequences. In contrast, ReMoS suffers from strong degradation, mainly because its spatio-temporal cross-attention mechanism requires very large memory. As a result, we had to reduce the batch size even when training on an NVIDIA H100 SXM5 GPU (80 GB HBM3 VRAM), which limited its performance. These results underline ReactionMamba’s robustness for stable and realistic long-term motion generation.
Lindy Hop (20 Frames $\sim$ 1s)
Model MPJPE MPJVE FIDFID $div_{gen}/div_{gt}$
ReMoS [20] 0.2194 0.0033 2.3097 1.0021
InterFormer [9] 0.1723 0.0046 2.7885 0.9936
Ready-to-React [66] (24 frames - 52 frames) 0.1483 0.0031 25.711 1.0032
ReactionMamba (Ours) 0.1635 0.0037 1.9852 1.0025
NTU120-AS (60 Frames $\sim$ 3s)
Model MPJPE MPJVE FIDFID $div_{gen}/div_{gt}$
ReMoS [20] 0.4577 0.0030 26.859 0.9654
InterFormer [9] 0.1865 0.0020 5.9255 1.0218
Ready-to-React [66] 0.5006 0.0037 12.292 0.9872
ReactionMamba (Ours) 0.0922 0.0020 0.0321 0.9816
TABLE II: Comparison of Results on long sequences generation. Bold denotes the best results, while underline indicates the second-best results.
InterX (200 Frames $\sim 10s$)
Model MPJPE MPJVE FIDFID $div_{gen}/div_{gt}$
ReMoS [20] 0.3318 0.0033 51.272 0.7891
InterFormer [9] 0.2394 0.0025 6.1582 0.9329
ReactionMamba (Ours) 0.0471 0.0007 0.0477 1.0026
Comparison of Inference Time. We present in table III the inference time needed to generate 3840 sequences of 60 frames each . We show the total time to generate all sequences in minutes, the time to generate a single sequence in seconds and the number of frames generated per second (FPS). We see that ReactionMamba is much faster than the state-of-the-art, specifically 900 times faster than Ready-to-React, 1000× faster than InterFormer, and more than 14,000× faster than ReMoS. These results underline ReactionMamba’s efficiency, showing that it not only scales to large datasets but also enables real-time long-horizon generation.
TABLE III: Inference Speed on NVIDIA Quadro RTX 6000 for 3840 NTU-AS120 Sequences of 60 Frames each
Model Total Time (min) Time per Sequence (s) FPS
ReMoS [20] 1691 26 2.3
Interformer [9] 122 1.9 32
Ready-to-React [66] 105 1.6 38
ReactionMamba (Ours) 0.117 0.0018 $\mathbf{33000}$
@@ V Results
Qualitative Results. We present in Figure 3 and Figure 4 visual results pertaining to long term retrogressive generation (1000 frames) on LindyHop and NTU120-AS push class respectively. Additional video results can be found in our supplementary material. In Figure 3 We observe that the qualitative results align well with the quantitative findings. The motion generated by ReactionMamba closely matches the ground truth and is comparable to Remos, while being hundreds of times faster. In contrast, Interformer’s motion loses coherence over time and fails to produce meaningful hand interactions. Similarly, Ready-to-React diverges after the first few hundred frames, with the reactor continuing its dance-like motion far away from the actor, and the generated hand motions becoming incoherent and poorly articulated. Meanwhile, in Figure 4 we see that ReactionMamba generate the best reaction to being pushed. Indeed, the other method show less intense motion as well as penetration between the two bodies. Overall, these results clearly demonstrate the superior quality, stability, and efficiency of ReactionMamba compared to existing methods.
@@
Figure 3: Visualization of sequence generated on Lindy Hop (1000 frames). In blue the action motion used as condition. In red the ground truth reaction and in other colors the reaction generated by the different models. See the supplementary material for their corresponding animations
Figure 4: Visualization of sequence generated on NTU120-AS push class. In blue the action motion used as condition. In red the ground truth reaction and in other colors the reaction generated by the different models. See the supplementary material for their corresponding animations
Ablation. We perform an ablation study to evaluate the influence of different components within our proposed ReactionMamba framework. We report ablation results in Table IV for LindyHop (20 frames).Five VAE variants are studied. S1 (VAE + Mamba + Concat Action + Concat Init-Pose) corresponds to our final model, where both the action stream and the initial pose are incorporated into the decoder via direct concatenation. S2 (S1 $\rightarrow$ Replace Mamba with Attention) keeps the same architecture as S1 but replaces all Mamba state space layers in the encoder and decoder with attention layers, which leads to significant degradation, especially in FID, confirming the effectiveness of Mamba for modeling long temporal dependencies. S3 (S1 $\rightarrow$ Remove Init-Pose Concatenation) removes the initial pose stream entirely in order to underline its necessity for generating consistent reaction motions. S4 (S1 $\rightarrow$ Replace Init-Pose Concatenation with Gating) introduces a dynamic gating mechanism, where the initial pose is injected into the latent sequence such that its influence gradually fades over time. While gating provides competitive performance, it consistently underperforms concatenation in reconstruction accuracy and stability, showing that directly concatenating the initial pose at every frame provides a stronger and more reliable conditioning signal. Finally, S5 (S1 $\rightarrow$ Replace Action Concatenation with Cross-Attention) replaces direct concatenation of the action stream with a cross-attention mechanism that enriches the latent sequence with action information. While cross-attention yields slightly improved distributional realism in some cases, it comes at the cost of higher reconstruction errors, suggesting that action concatenation remains the more balanced and robust strategy overall. To conclude, for short sequences, concatenation clearly dominates (S1). For long sequences, gating or cross-attention can optimize a single metric (MPJPE or FID), but S1 offers the best overall balance of accuracy (MPJPE/MPJVE), distribution realism (FID), and stable diversity close to 1.0, while keeping the model simple, controllable, and robust to training instabilities. We therefore adopt S1 (VAE + Mamba + Concat(Action) + Concat(Init-Pose)) as our final architecture, and report S4/S5 as targeted alternatives when one metric (MPJPE or FID) is prioritized over the global trade-off.
Scenario MPJPE ↓ MPJVE ↓ FID ↓ $div_{gen}/div_{gt}$
S1 0.1634 0.00374 1.9852 1.0025
S2 0.1833 0.00407 2.2392 1.0037
S3 0.2299 0.00543 2.3857 1.0157
S4 0.1773 0.00410 2.0318 0.9836
S5 0.2067 0.00427 2.0915 0.9925
TABLE IV: Ablation study on ReactionMamba. Results for LindyHop sequences of 20 frames each
VI Limitations and future work
VII Conclusion
Acknowledgements

### v2

@@ XI Metric Definitions
A lower FID indicates that the generated data distribution closely matches the real data distribution, reflecting higher quality and diversity in the generated motions  [10, 19]. For all methods, FID is computed using the same transformer based motion encoder trained on the corresponding dataset and joint representation.
@@
where $S^{i}$ and $S^{j}$ are two sampled motion sequences from set $S$ (real or generated), and $P$ is the set of unique pairs randomly selected. A diversity score close to the ground truth diversity score reflects greater variability within the motion set and preservation of the original diversity.
XII More visual results
@@
We illustrate long-term retrogressive generation on LindyHop (1000 frames) in Figure 5. We observe that the qualitative results align well with the quantitative findings. The motion generated by ReactionMamba closely matches the ground truth and is comparable to Remos, while being hundreds of times faster. Interformer and the Seq2SeqMamba fails to produce meaningful hand interactions. Similarly, Ready-to-React diverges after the first few hundred frames, with the reactor continuing its dance-like motion far away from the actor, and the generated hand motions becoming incoherent and poorly articulated. Meanwhile, Overall, these results clearly demonstrate the better quality, stability, and efficiency of ReactionMamba compared to existing methods.
@@
Figure 5: Visualization of sequence generated on Lindy Hop dance. In blue the action motion used as condition. In green the ground truth reaction and in other colors the reaction generated by the different models. See the supplementary material for their corresponding animations